A teacher allows the students to each throw out their lowest test score in order to increase their overall grade. The original mean score of 5 tests is 75.2 After the removal of the lowest test, the mean score becomes 80.25. What is the score of the removal test?

Answer by stanbon(75887)

A teacher allows the students to each throw out their lowest test score in order to increase their overall grade. The original mean score of 5 tests is 75.2 After the removal of the lowest test, the mean score becomes 80.25. What is the score of the removal test?
----------------------------------------
Original test score sum:
sum = 5*75.2 = 376 points
-----------------------------------
Let "x" be the score that was removed.
Points remaining = 376-x
------------
Equation:
New average = (376-x)/4 = 80.25
376-x = 4*80.25
276-x = 321
x = -45
------------
The score thrown out was 45
---------------
Cheers,
Stan H.